Approximately how many times louder is a 145-dB sound than a 85-dub sound?


1
Expert's answer
2021-10-05T10:07:21-0400

For the 145-dB-sound:


L1=10log10P1P0 P1=10L1/10P0,L_1=10\log_{10}\frac{P_1}{P_0}\space\rightarrow P_1=10^{L_1/10}P_0,

for the 85-dB-sound:


L2=10log10P2P0 P2=10L2/10P0.L_2=10\log_{10}\frac{P_2}{P_0}\space\rightarrow P_2=10^{L_2/10}P_0.

Find the ratio:


P1P2=10(L1/10L2/10)=106.\frac{P_1}{P_2}=10^{(L_1/10-L_2/10)}=10^6.